The width of a rectangle is one-fifth the length. The perimeter of the rectangle is 30 feet. What is the width of the rectangle?
how tdo you do this ?

Answer by prateekagrawal(56)

let length is 5x.
then width= 5x/5 = x
.
so, perimeter=2(length+width)
30=2(5x+x)
30= 2*6x
30=12x
x=30/12 = 2.5 feet
so, length= 5*2.5 = 12.5 feet
and width= 2.5 feet